# From Bias to Balance: Exploring and Mitigating Spatial Bias in LVLMs

### v1

@@ 4.2 Is Vision Encoder Understanding Spatially Robust?
Task Design. Inspired by Radford et al. (2021), we determine whether the vision encoder can maintain consistent semantic understanding when key image locations change by analyzing the similarity between the visual features input to the LLMs and the embeddings of the corresponding caption. Specifically, we first randomly select 1,000 key image-caption pairs $(I^{\prime}_{m},C^{\prime}_{m})$ from LAION dataset. As illustrated in Figure 8(b), each image $I^{\prime}_{m}$ is then pasted onto a white background at every position $n$, serving as synthetic image inputs $\bm{I^{\prime}}_{m,n}$ for the LVLMs. We measure the semantic understanding capability of the vision encoder by evaluating the cosine similarity between the embedding representations of caption $C^{\prime}_{m}$ and visual features of $\bm{I^{\prime}}_{m,n}$ input to LLM.
$\text{Similarity}=\cos(g(f_{v}(\bm{I^{\prime}}_{m,n})),E(C^{\prime}_{m})),$ (1)
@@
Figure 3: Results on understanding abilities across each Position $n$. Each cell reports the similarity between visual features and text embeddings input to the LLMs when the key image $I_{m}$ is placed at a specific grid location $n$ in the $3\times 3$ composite image $\bm{I}_{m,n}$.
@@
Results and Findings. As shown in Figure 3, the similarity scores remain highly stable regardless of the spatial placement of the key image across all five models. This consistency indicates that the vision encoder reliably extracts semantically aligned representations of the key image, independent of its spatial location in the composite. Moreover, combining the results in Figure 1, we find a trend that LVLMs achieving higher similarity scores (e.g., Gemma3 and Qwen2.5-VL) also exhibit stronger and more stable performance on our probing tasks, whereas models with lower similarity scores (e.g., LLaVA-v1.6 and LLaVA-NeXT) perform worse and show greater sensitivity to positional changes. This correlation further validates that robust semantic alignment between visual features and textual embeddings is a key factor behind reliable multimodal understanding.
Figure 4: Overall view of inference in LVLMs with traditional position assignment and the proposed balanced position assignment.
5 Balanced Position Assignment (BaPA)
Based on the above observation, we assume that the root of spatial bias may be traced to the imbalance inherent in the position embeddings within the LLM component of LVLMs.
5.1 The Baseline Rotary Position Embedding
Currently, most LVLMs employ Rotary Position Embedding (RoPE) (Su et al., 2024) as the position encoding method for the LLM component. Unlike absolute position embeddings that add a fixed vector to each token representation, RoPE encodes relative positional information directly into the attention mechanism through rotation in a complex plane. Specifically, for a token embedding $x_{p}$ at position $p$, RoPE applies a rotary matrix $\bm{R}_{\Theta,p}^{d}$ with pre-defined parameters $\Theta=\{\theta_{i}=10000^{-2(i-1)/d},i\in[1,2,\dots,d/2]\}$. Thus, during the self-attention computation (Vaswani et al., 2017), the dot product between query and key at position $p$ and $q$ becomes
$q_{p}^{\mathsf{T}}k_{q}=(\bm{R}_{\Theta,p}^{d}W_{q}x_{p})^{\mathsf{T}}(\bm{R}_{\Theta,q}^{d}W_{k}x_{q})=x_{p}^{\mathsf{T}}W_{q}\bm{R}_{\Theta,{q-p}}^{d}W_{k}x_{q}$ (2)
where $\bm{R}_{\Theta,{q-p}}^{d}=(\bm{R}_{\Theta,p}^{d})^{\mathsf{T}}\bm{R}_{\Theta,q}^{d}$ and qpq-p represents the relative position between $q_{p}$ and $k_{q}$. According to Equation 2, RoPE naturally introduces imbalance, where the result of the inner product, i.e., interaction strength, is influenced by the relative distance between tokens. While this imbalance is well-suited for capturing autoregressive dependencies in text, we posit that it is ill-suited for facilitating effective cross-modal interactions below.
@@ 5.2 The Proposed Method
In traditional LVLMs, visual features extracted from the vision encoder are typically arranged into a 1-D sequential order following a raster-scan strategy as shown in Figure 4(a). Due to the imbalance property of RoPE (Peng et al., 2024), image tokens at different positions in the sequence have unequal influence on the same text token during cross-modal interaction. However, intuitively, since the vision encoder has already modeled local pixel relationships and global spatial structures, image tokens should be treated as semantically equivalent from the LLM’s perspective.
To overcome this imbalance, we propose Balanced Position Assignment (BaPA), a simple yet effective modification to the position assignment of image tokens. In particular, as shown in Figure 4(b), for a a multimodal sequence $Z=\{s_{1},\dots,s_{i},\tilde{v}_{1},\dots,\tilde{v}_{j},x_{1},\dots,x_{k}\}$, where $\{\tilde{v}_{1},\dots,\tilde{v}_{j}\}$ denotes the image tokens derived from vision encoder and projector, $\{s_{1},\dots,s_{i}\}$ and $\{x_{1},\dots,x_{k}\}$ denotes the embeddings of system prompt and user prompt, respectively, instead of endowing each image token $\tilde{v}_{k}$ with distinct position assignment $p_{\tilde{v}_{k}}$, BaPA enforces
$p_{\tilde{v}_{1}}=p_{\tilde{v}_{2}}=\dots=p_{\tilde{v}_{j}}=p_{img}=i$ (3)
where $i$ and $j$ are the number of the system prompt tokens and image tokens, respectively. Here, in order to ensure the continuity of the overall position assignment, we set $p_{img}$ to the value of the original $p_{\tilde{v}_{1}}$, i.e., $i$. Therefore, the final position assignment for the entire input of the LLM backbone is as follows:
$\text{Position\_ids}=\{\underbrace{0,1,\dots,i-1}_{\text{system prompt}},\underbrace{i,i,\dots,i}_{\text{image tokens}},\underbrace{i+1,i+2,\dots,i+k-1}_{\text{user prompt}}\}$ (4)
This modification removes the artificial positional imbalance introduced by RoPE or similar encodings, ensuring that each image token has equal importance when interacting with textual tokens.
Table 1: Results on our probe dataset where 0-8 denotes the different grid positions of key image in the composite images, AvgAvg and $\Delta$ denote the average accuracy and variance across all positions, respectively. The results with an increase after BaPA are bolded. Noted that the results of BaPA are obtained without training the baseline LVLMs.
Position 0 1 2 3 4 5 6 7 8 Avg $\uparrow$ $\Delta$ $\downarrow$
Gemma3 83.77 83.05 78.11 85.11 84.43 78.98 84.41 82.87 77.60 0.82  8.75
Gemma3-BaPA 93.03 93.10 91.05 92.66 92.97 90.01 92.49 92.94 90.48 92.08  1.49
LLaVA-NeXT 68.91 77.62 70.89 58.54 69.82 62.52 55.20 63.89 57.15 64.95 54.90
LLaVA-NeXT-BaPA 96.44 96.35 96.52 96.31 96.12 96.47 96.42 96.46 96.63 96.41  0.02
LLaVA-v1.6-BaPA 75.71 77.75 71.67 70.98 65.87 68.35 65.87 72.40 68.81 70.82 16.79
LLaVA-v1.6-BaPA 94.89 84.47 93.63 94.18 93.79 92.71 94.05 93.77 92.55 93.78  9.97
Qwen2.5-VL-7B 81.86 80.00 81.67 83.81 81.71 83.58 84.15 82.62 83.05 82.49  1.74
Qwen2.5-VL-7B-BaPA 80.95 78.71 79.92 81.95 80.43 82.23 83.41 81.54 82.38 81.28  2.06
Qwen2.5-VL-32B 82.25 81.82 82.58 82.39 81.47 81.66 82.52 81.92 81.00 81.96  0.28
Qwen2.5-VL-32B-BaPA 82.31 81.69 82.31 82.72 81.66 81.53 84.18 83.57 82.34 82.48  0.80
5.3 Experiments on Probe Dataset
Experimental Settings. We first evaluate the effectiveness of the proposed BaPA on our probe dataset. Since the training data of mainstream LVLMs is not publicly available, retraining these models with a modified position embedding is infeasible. Thus, we directly apply BaPA to the inference stage of existing LVLMs without any retraining, making our approach lightweight and broadly applicable. We test the same five representative LVLMs under a zero-shot setting.
Results and Findings. Table 1 reports the performance of each LVLM and their BaPA-enhanced variants on our probe dataset, where the key image is placed at different grid positions (0–8). For each model, we further report the average accuracy (Avg) and variance ($\Delta$) across all positions. The results highlight three major findings.
  • For average accuracy, applying BaPA leads to a substantial improvement across almost all models, with particularly remarkable gains observed for previous underperforming models like LLaVA-NeXT and LLaVA-v1.6. This suggests that the baseline’s unbalanced position embeddings hinder the full utilization of visual information, and rectifying this imbalance enables LVLMs to exploit visual features more effectively for cross-modal interaction.
  • For spatial robustness, BaPA effectively reduces the accuracy variance across different positions. Notably, LLaVA-NeXT, which originally suffered from severe instability ($\Delta=54.90$), achieves near-uniform performance ($\Delta=0.02$) after applying BaPA. This confirms that the root of spatial bias lies in the unbalanced position assignment within LLMs, and that equalizing positional importance directly mitigates this inconsistency.
  • For model scale, as the model parameters increase from 7B to 32B, BaPA delivers a more pronounced overall accuracy improvement for Qwen2.5-VL, rising from 81.28 to 82.48. This suggests that the proposed method is effective in enhancing the performance of larger models. Moreover, owing to its improved MRoPE (Bai et al., 2025), the spatial bias inherent in Qwen2.5-VL is less significant, hence our BaPA provides no improvement on variance.
5.4 Experiments on Downstream Tasks
Experimental Settings.
To further validate the effectiveness of BaPA beyond the probing task, we evaluate its performance on general downstream multimodal tasks with three representative LVLMs, including Gemma3-8B (Team et al., 2025), LLaVA-v1.6-Mistral-7B (Liu et al., 2024), Qwen2.5-VL-7B (Bai et al., 2025). Considering that downstream tasks involve more complex image–text interactions than our probe dataset, we perform lightweight LoRA (Hu et al., 2022) fine-tuning with 10K instruct-tuning data from LLaVA to adapt these models to the new position encoding scheme. We additionally train LLaVA-v1.5 (Liu et al., 2024) with BaPA from scratch to compare with CCA, which is a positional alignment strategy proposed by Xing et al. (2024) to address the object hallucination of LVLMs. Following (Liu et al., 2024), we only retained samples that contain exactly one image in the input for each dataset during evaluation. More details are available in Appendix C.3.
Table 2: Results on downstream tasks. The results with an increase after BaPA are bolded.
@@ 5.4 Experiments on Downstream Tasks
Models $\text{MMMU-Pro}_{\text{direct}}$ $\text{ScienceQA}_{\text{image}}$ $\text{CRPE}_{\text{relation}}$ HallusionBench
@@
Gemma3 0.4357 0.2536 0.8156 0.7002 0.6393
Gemma3-BaPA 0.4162 0.2944 0.8180 0.6819 0.6015
Qwen2.5-VL 0.4763 0.3427 0.7898 0.7658 0.7066
Qwen2.5-VL-BaPA 0.4915 0.3735 0.8909 0.7690 0.6909
LLaVA-v1.6 0.3390 0.2015 0.7288 0.6884 0.5363
LLaVA-v1.6-BaPA 0.3365 0.1927 0.7288 0.6905 0.5489
LLaVA-v1.5 0.3522 0.1908 0.6951 0.6671 0.4825
LLaVA-v1.5-CCA 0.3340 0.1864 0.6966 0.6581 0.4921
LLaVA-v1.5-BaPA 0.3440 0.1952 0.7010 0.6760 0.5205
MMMU-Pro. We first evaluate the proposed method on MMMU-Pro (Yue et al., 2025), which is an enhanced multimodal benchmark designed to rigorously assess the true understanding capabilities of LVLMs. As shown in Table 2, BaPA yields mixed effects on such complex reasoning tasks, while CCA typically reduces the performance of LVLMs. BaPA provides a slight decrease for Gemma3 in the 4-option setting, yet it confers a significant advantage (from 0.2536 to 0.2775) in the more challenging 10-option setting. For Qwen2.5-VL, BaPA consistently enhances performance in both settings, especially for 10-option with 3% accuracy improvement. This suggests that the benefits of BaPA are most pronounced in tasks that require richer contextual reasoning from multimodal inputs.
@@
ScienceQA. We then evaluate our method on the ScienceQA dataset (Lu et al., 2022), which is collected from elementary and high school science curricula. The results in Table 2 show that BaPA can consistently improve or maintain LVLMs’ performance on science-related tasks. The most notable gain is observed in Qwen2.5-VL with a substantial improvement of 10%. This suggests that BaPA effectively reduces positional imbalance, leading to better integration of visual and textual information for precise knowledge-based reasoning.
@@
Hallucination Benchmark. We also evaluate BaPA on two comprehensive multimodal hallucination benchmarks, i.e., CRPE (Wang et al., 2024b) and HallusionBench (Guan et al., 2024). According to Table 2, we find that BaPA generally reduces hallucination tendencies and improves the prediction reliability of LVLMs. For instance, BaPA improves the performance of LLaVA-v1.6 on HallusionBench from 0.5363 to 0.5489, while Qwen2.5-VL achieves a small gain on CRPE. Although some models, such as Gemma3, exhibit minor drops, the overall trend suggests that BaPA helps alleviate the negative effects of spatial bias, leading to more stable and trustworthy predictions.
Table 3: Results on MME benchmark. The results with an increase after BaPA are bolded.
Model Object-level Attribute-level Total
existence count position color
Gemma3 190.00 165.00 98.33 158.33 611.66
Gemma3-BaPA 200.00 151.67 111.67 153.33 616.67
Qwen2.5-VL 200.00 155.00 160.00 195.00 710.00
Qwen2.5-VL-BaPA 200.00 163.33 175.00 195.00 733.33
LLaVA-v1.6 200.00 155.00 133.33 185.00 673.33
LLaVA-v1.6-BaPA 200.00 153.33 145.00 175.00 673.33
LLaVA-v1.5 175.67 124.67 114.00 151.00 565.33
LLaVA-v1.5-UAC 190.00 155.00 128.33 165.00 638.33
LLaVA-v1.5-CCA 190.00 148.33 128.33 175.00 641.66
LLaVA-v1.5-BaPA 190.00 153.33 135.00 170.00 648.33
MME. The MME benchmark (Fu et al., 2023) is a comprehensive evaluation benchmark for LVLMs. Following Xing et al. (2024), we evaluate our method on four perception subtasks that assess object-level and attribute-level hallucinations. Here we also compare the results of UAC (Zhu et al., 2025c) [^2] Here we compare with UAC rather than DAC proposed by Zhu et al. (2025c), as DAC is specifically fine-tuned on the 5.4k data that is constructed with ground truth object labels, making a direct comparison unfair., a training-free attention calibration method for object hallucination. Due to the differing training strategies of the two baselines, we employ a compromise approach for comparison, i.e., fine-tuning LLaVA-v1.5 using the BaPA with LoRA on 10K samples as other models. The results in Table 3 show that BaPA consistently preserves or improves model performance on the MME benchmark, especially at the attribute level, where position imbalance has the strongest negative effect. Compared with CCA and UAC, our BaPA achieves better results on most tasks and total scores despite being a much simpler positional adjustment. Overall, these findings indicate that equalizing image-token position embeddings is a compact, effective way to strengthen LVLMs’ grounding and reduce hallucination without heavy re-training or complex architectural changes.

### v2

@@ 4.2 Is Vision Encoder Understanding Spatially Robust?
Task Design. Inspired by Radford et al. (2021), we determine whether the vision encoder can maintain consistent semantic understanding when key image locations change by analyzing the similarity between the visual features input to the LLMs and the embeddings of the corresponding caption. Specifically, we first randomly select 1,000 key image-caption pairs $(I^{\prime}_{m},C^{\prime}_{m})$ from LAION dataset. As illustrated in Figure 7(b), each image $I^{\prime}_{m}$ is then pasted onto a white background at every position $n$, serving as synthetic image inputs $\boldsymbol{I^{\prime}}_{m,n}$ for the LVLMs. We measure the semantic understanding capability of the vision encoder by evaluating the cosine similarity between the embedding representations of caption $C^{\prime}_{m}$ and visual features of $\boldsymbol{I^{\prime}}_{m,n}$ input to LLM.
$\text{Similarity}=\cos(g(f_{v}(\boldsymbol{I^{\prime}}_{m,n})),E(C^{\prime}_{m})),$ (1)
@@
Results and Findings. As shown in Figure 3, the similarity scores remain highly stable regardless of the spatial placement of the key image across all five models. This consistency indicates that the vision encoder reliably extracts semantically aligned representations of the key image, independent of its spatial location in the composite. Moreover, combining the results in Figure 1, we find a trend that LVLMs achieving higher similarity scores (e.g., Qwen2.5-VL, Gemma3n and InternVL3) also exhibit stronger and more stable performance on our probing tasks, whereas models with lower similarity scores (e.g., LLaVA-v1.6) perform worse and show greater sensitivity to positional changes. This correlation further validates that robust semantic alignment between visual features and textual embeddings is a key factor behind reliable multimodal understanding.
5 Adaptive Global Context Injection (AGCI)
Based on the above observation, we assume that the root of spatial bias may be traced to the cross-modal reasoning within the LLM backbone, where image tokens are processed under causal attention constraints. This mechanism is incompatible with the attention in the visual encoder, weakening the availability of global visual semantics during cross-modal reasoning.
5.1 Attention Mechanism
Attention is a core operation in both vision and language models, which dynamically allocate weights to different parts of an input sequence, allowing the model to focus on the most relevant features for a given context. Formally,
$\text{Attention}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\text{softmax}\!\left(\frac{\boldsymbol{Q}\boldsymbol{K}^{\mathsf{T}}}{\sqrt{d_{k}}}+\boldsymbol{M}\right)\boldsymbol{V},$ (2)
where $\boldsymbol{Q}=\text{Pos}(\boldsymbol{W}_{q}\boldsymbol{X},\boldsymbol{PE})$, $\boldsymbol{K}=\text{Pos}(\boldsymbol{W}_{k}\boldsymbol{X},\boldsymbol{PE})$, and $\boldsymbol{V}=\text{Pos}(\boldsymbol{W}_{v}\boldsymbol{X},\boldsymbol{PE})$ denotes the query, key and value matrices derived from the input $\boldsymbol{X}$ and the position encoding matrix $\boldsymbol{PE}$. $\boldsymbol{M}$ is the mask matrix where elements are typically set to -$\infty$ or 0 to indicate the visibility between tokens.
Current LVLMs typically adopt a Vision Transformer (ViT) (Dosovitskiy et al., 2021) as the visual encoder. It extracts a sequence of visual tokens (patch embeddings) and applies bidirectional self-attention (i.e., each token can attend to all other tokens in the same image), enabling global contextual aggregation of visual information. Formally, the vision encoder uses a full attention mask
$\boldsymbol{M}_{\text{bi}}=\mathbf{0}_{L\times L}$ (3)
where $L$ is the number of visual tokens.
In contrast, the LLM backbone uses causal self-attention, where each token can only attend to itself and previous tokens. The corresponding mask is
$\boldsymbol{M}_{\text{causal}}[i,j]=\begin{cases}0,&j\leq i,\\ -\infty,&j>i,\end{cases}$ (4)
which enforces a strictly lower-triangular visibility pattern.
Table 1: Results on our probe dataset where 0-8 denotes the different grid positions of key image in the composite images, AvgAvg and $\Delta$ denote the average accuracy and variance across all positions, respectively. The results with an increase after AGCI are bolded.
Position 0 1 2 3 4 5 6 7 8 Avg $\uparrow$ $\Delta$ $\downarrow$
Gemma3n-E4B 90.08 91.36 89.38 88.75 91.62 89.25 86.75 89.53 87.19 89.32 2.39
Gemma3n-E4B-AGCI 97.64 97.76 97.19 97.33 97.63 96.73 96.49 97.11 96.30 97.13 0.25
LLaVA-v1.6-7B 75.71 77.75 71.67 70.98 65.87 68.35 65.87 72.40 68.81 70.82 14.93
LLaVA-v1.6-7B-AGCI 84.84 85.08 80.90 80.52 76.86 77.77 79.19 83.07 78.49 80.75 8.05
InternVL3-8B 83.59 82.42 86.96 83.89 81.79 84.70 86.41 82.37 83.67 83.98 2.82
InternVL3-8B-AGCI 85.72 83.58 87.79 86.05 82.94 85.88 87.89 84.35 85.32 85.50 2.56
Qwen3-VL-8B 73.79 74.55 76.24 70.24 71.78 73.96 67.97 70.28 73.39 72.47 5.98
Qwen3-VL-8B-AGCI 79.50 80.01 81.13 76.88 77.68 79.55 75.44 76.09 79.64 87.44 3.45
Qwen2.5-VL-7B 81.86 80.00 81.67 83.81 81.71 83.58 84.15 82.62 83.05 82.49 1.54
Qwen2.5-VL-7B-AGCI 89.25 87.98 89.29 90.20 88.30 90.06 91.04 89.73 90.23 89.56 0.84
Qwen2.5-VL-32B 82.25 81.82 82.58 82.39 81.47 81.66 82.52 81.92 81.00 81.96 0.25
Qwen2.5-VL-32B-AGCI 83.58 83.23 83.69 83.48 82.67 82.68 83.62 83.20 81.93 83.12 0.31
Qwen2.5-VL-72B 70.33 70.61 71.64 69.93 70.39 70.84 68.60 68.40 70.36 70.12 0.95
Qwen2.5-VL-72B-AGCI 86.11 85.98 86.62 85.88 85.73 86.00 86.08 85.80 86.79 86.11 0.12
@@ 5.2 The Proposed Method
In traditional LVLMs, the unidirectional constraint in LLM differs fundamentally from the bidirectional attention within vison encoder. It may hinder global context exchange among image tokens during cross-modal processing, causing image tokens to gradually lose access to holistic image information as they are propagated through the Transformer layers.
To mitigate this issue, we propose Adaptive Global Context Injection (AGCI), which injects shared global image context into each image token in an adaptive manner. Given the image token sequence $\{\tilde{v}_{i}\}_{i=1}^{L}$, we first compute the mean representation as global image context
$\tilde{g}=\frac{1}{L}\sum_{i=1}^{L}\tilde{v}_{i}.$ (5)
Then, for each image token $\tilde{v}_{i}$, we measure its semantic similarity to the global context by
[…]
Finally, the global context is injected into each image token with an adaptive strength proportional to $(1-w_{i})$:
$\tilde{v}_{i}^{\prime}=\tilde{v}_{i}+\lambda(1-w_{i})\,\tilde{g},\qquad i=1,\dots,L.$ (7)
where $\lambda$ is a hyperparameter to cap the maximum amount of injected global information, serving as a safety valve to better preserve fine-grained visual details. This design ensures the tokens with lower global semantic relevance receive stronger global-context supplementation, helping preserve holistic image information during cross-modal interaction. We provide a theoretical analysis of AGCI from the perspective of mutual information in Appendix C.
Table 2: Results on downstream tasks. The best results are bolded.
@@
Models $\text{MMMU-Pro}_{\text{direct}}$ $\text{ScienceQA}_{\text{image}}$ $\text{CRPE}_{\text{relation}}$ HallusionBench TextVQA
@@
Gemma3n 36.53 23.85 83.94 69.22 66.04 68.97
Gemma3n-AGCI 41.48 28.37 83.59 69.97 63.04 70.54
LLaVA-v1.6 33.90 20.15 72.88 68.84 53.63 63.85
LLaVA-v1.6-AGCI 34.21 20.09 73.33 68.86 54.47 65.68
InternVL3 53.17 37.92 98.22 76.15 65.09 81.12
InternVL3-AGCI 53.35 37.60 98.41 75.63 66.14 81.81
Qwen3-VL 38.42 27.93 94.79 77.68 73.50 80.34
Qwen3-VL-AGCI 45.45 34.78 94.05 77.92 72.56 81.13
Qwen2.5-VL 41.18 28.81 88.60 76.67 71.08 77.72
Qwen2.5-VL-AGCI 44.57 31.64 88.84 76.95 69.40 82.65
6 Experiments
6.1 Experimental Settings
In this section, we evaluate the effectiveness of the proposed AGCI on our probe dataset and six downstream multimodal tasks. Since the training data of mainstream LVLMs is not publicly available, retraining these models with AGCI is infeasible. Thus, we perform lightweight LoRA (Hu et al., 2022) fine-tuning with 10K instruct-tuning data from LLaVA (Liu et al., 2023) to adapt each model to the modified image token representation. For LLaVA-v1.6, we set $\lambda$ to $0.3$, while all other LVLMs are set to $0.5$. Following (Liu et al., 2024a), we only retained samples that contain exactly one image in the input for each dataset during the evaluation on downstream tasks. We test the same five representative LVLMs as above under a zero-shot setting. More details are available in Appendix D.2.
6.2 Results on Probing Task
Table 1 reports the performance of each LVLM and their AGCI-enhanced variants on our probing task, where the key image is placed at different grid positions (0–8). For each model, we further report the average accuracy (Avg) and variance ($\Delta$) across all positions. The results highlight three major findings.
For average accuracy, AGCI consistently improves performance across all evaluated LVLMs, with particularly large gains upon previously underperforming models (e.g., LLaVA-v1.6 and Qwen3-VL). This suggests that the asymmetric interaction of image tokens caused by causal attention hinders the full utilization of visual information, and injecting shared global visual context improves the semantic accessibility of image tokens and strengthens cross-modal evidence aggregation.
For spatial robustness, AGCI effectively reduces the accuracy variance across different positions for almost all models, indicating improved invariance to spatial shifts of critical visual information. Notably, LLaVA-v1.6, which originally suffers from severe instability, exhibits substantial variance reductions after applying AGCI. This confirms that the cause of spatial bias lies in the disruption of global information flow within the LLM backbone, and preserving such information effectively alleviates this inconsistency.
For model scale, as the model parameters increase from 7B to 72B, AGCI delivers a more pronounced overall accuracy improvement while maintaining low variance, particularly for Qwen2.5-VL-72B, where accuracy surges from 70.12 to 86.11, with variance also dropping from 0.95 to 0.12. This indicates that spatial bias is not merely a small-model artifact, and explicitly reinforcing global context can further stabilize cross-position performance even for larger LVLMs.
6.3 Results on Downstream Tasks
MMMU-Pro. We first evaluate the proposed method on MMMU-Pro (Yue et al., 2025), which is an enhanced multimodal benchmark designed to rigorously assess the true understanding capabilities of LVLMs. As shown in Table 2, AGCI improves the performance of almost all models when confronted with such complex reasoning tasks, with particularly notable gains on both settings. This indicates that mitigating spatial bias benefits complex reasoning scenarios where models must integrate multiple visual cues and maintain a coherent global understanding of the image. The larger improvements observed on the 10-option setting further suggest that AGCI enhances the robustness of fine-grained reasoning under increased decision complexity.
@@ 6.3 Results on Downstream Tasks
ScienceQA. We then evaluate our method on the ScienceQA dataset (Lu et al., 2022), which is collected from elementary and high school science curricula. The results in Table 2 show that AGCI can consistently improve or maintain LVLMs’ performance on science-related tasks. These results suggest that injecting global visual semantics helps the LVLM better align visual evidence with textual knowledge, leading to better integration of visual and textual information for precise knowledge-based reasoning.
@@
Hallucination Benchmark. We also evaluate AGCI on two comprehensive multimodal hallucination benchmarks, i.e., CRPE (Wang et al., 2024b) and HallusionBench (Guan et al., 2024). According to Table 2, we find that AGCI generally reduces hallucination tendencies and improves the prediction reliability of most LVLMs. For instance, AGCI improves the performance of InternVL3 on HallusionBench from 65.09 to 66.14, while other models achieve a small gain on CRPE. Although some models, such as Gemma3n, exhibit minor drops, the overall trend suggests that AGCI helps alleviate the negative effects of spatial bias, leading to more stable and trustworthy predictions.
TextVQA. TextVQA (Singh et al., 2019) evaluates LVLM’s ability to recognize and reason over scene text, which is highly sensitive to spatial localization. Hence, we further validate the effectiveness of our method on this task. Based on Table 2, we can observe that AGCI consistently improves performance across all models, with especially strong gains for Qwen2.5-VL (from 77.72 to 82.65) and LLaVA-v1.6 (+4.93). This suggests that restoring global visual semantics helps the LLM better aggregate dispersed textual regions in the image, alleviating spatial bias and improving the alignment between textual queries and relevant OCR tokens.
Table 3: Accuracy on POPE MSCOCO dataset, Avg. indicates the average accuracy. The best results are bolded. The results of baselines based on LLaVA-v1.5 are reported by Xing et al. (2024). The complete results are presented in Table 6.
Models Adv. Pop. Rand. Avg.
Gemma3n 86.77 87.90 88.93 87.87
Gemma3n-AGCI 87.07 88.37 89.40 88.28
LLaVA-v1.6 87.10 88.70 90.80 88.87
LLaVA-v1.6-AGCI 87.17 89.23 91.40 89.27
InternVL3 86.43 88.00 89.20 87.88
InternVL3-AGCI 86.57 88.13 89.57 88.09
Qwen3-VL 82.89 84.10 86.54 84.51
Qwen3-VL-AGCI 84.67 86.57 89.63 86.96
Qwen2.5-VL 87.77 90.67 94.13 90.86
Qwen2.5-VL-AGCI 89.00 90.73 93.17 90.97
LLaVA-v1.5 78.96 81.88 83.29 81.38
LLaVA-v1.5-VCD 80.88 85.38 87.73 84.66
LLaVA-v1.5-RLHF 82.60 83.90 85.90 84.13
LLaVA-v1.5-CCA 85.67 86.87 88.03 86.86
LLaVA-v1.5-AGCI 85.50 86.87 88.57 86.98
POPE. POPE (Li et al., 2023b) provides a detailed evaluation of object hallucination in LVLMs. To demonstrate the benefits of AGCI, we evaluate our method with POPE built on the validation set of MSCOCO (Lin et al., 2014) under adversarial (Adv.), popular (Pop.), and random (Rand.) negative sampling strategies. We additionally train LLaVA-v1.5 (Liu et al., 2024a) with AGCI from scratch to compare three baseline methods: VCD (Leng et al., 2024), RLHF (Sun et al., 2024), and CCA (Xing et al., 2024). As shown in Table 3, AGCI consistently improves performance on POPE across different negative sampling strategies, indicating enhanced visual grounding and reduced hallucination. Notably, Qwen3-VL benefits substantially from AGCI, with its average accuracy improving from 84.51 to 86.96, showing stronger robustness under adversarial and popular sampling. Similar improvements are observed for LLaVA-v1.6 and InternVL3, both achieving higher average accuracy after applying AGCI. Meanwhile, for LLaVA-v1.5, our AGCI still consistently outperforms baselines across most evaluation settings, even though our method is not specifically designed for the object hallucination task. These results suggest that injecting global visual semantics effectively reduces reliance on language priors and improves model’s robustness across diverse LVLM architectures.
Figure 4: An example of visualization of information flow from image token to text token. More cases are provided in Appendix E.